OpenStudy (anonymous):

what is 100^2-99^2+98^2-97^2+96^2-95^2+....+4^2-3^2+2^2-1^2? I need to find an expression where n is the largest number being squared. I notice that every odd number is minus and even number is a positive.

OpenStudy (mathmath333):

\(\large\tt \color{black}{100^2-99^2+98^2-97^2+96^2-95^2+....+4^2-3^2+2^2-1^2}\) if u write it in reverse it looks \(\large\tt \color{black}{2^2-1^2+4^2-3^2...........100^2-99^2}\) that is \(\large\tt \color{black}{3+7+11............+199}\)

OpenStudy (mathmath333):

\(\large\tt \color{black}{sum=50\times\dfrac{199+3}{2}=5050}\)

OpenStudy (mathmath333):

where \(\large\tt \color{black}{50}\) is the number of terms in the series